A student wearing frictionless in-line skates in a horizontal surface is pushed by a friend with a constant force of 45N. How far must the student be pushed, starting from rest her final kinetic energy is 352J?
1
Expert's answer
2013-01-08T11:43:42-0500
T=A=F*d
T - the kinetic energy
A - the work done by a constant force of magnitude F on a body that moves a distance d in the direction of the force is the product
d=T/F=325/45=7.22 m